OpenStudy (anonymous):

DANGER! need help! please? please help! crashing for finals! find cos(u+v) given that sin u=3/5 and cos v=-5/13 (both u and v are in Quad 2

OpenStudy (anonymous):

please help?

OpenStudy (anonymous):

\[ \cos(u+v) = \cos(u) \cos(v) - \sin(u) \sin(v) \]

OpenStudy (anonymous):

\[ \cos^2(u)= 1 - \frac 9{25}= \frac {16}{25} \]

OpenStudy (anonymous):

\[ \sin^2(v)= 1 - \frac {25}{169}= \frac {144}{169} \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

i having a little trouble can u explain what u just did please? step-by-step?

OpenStudy (anonymous):

Be careful u and v are in quadrant 2

OpenStudy (anonymous):

I used \[ \cos^2(x) + \sin^2(x) =1 \]

OpenStudy (anonymous):

okay so u used this formula and plugged in the values?

OpenStudy (anonymous):

that were given?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so it'll be cos^2(x)+sin^2(x)=1 and the i plug in the given info so it''l be cos^2(-5/13)+sin^2(3/5)=1?

OpenStudy (anonymous):

you should obtain \[ \cos(u) = -\frac 4 5\\ \sin(u) =\frac {12}{13} \]

OpenStudy (anonymous):

how?